# On the contraction properties of Sinkhorn semigroups

### v2

@@ D.3 Lyapunov inequalities
$e^{\nu_{V}(V)}~e^{-({\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}\leq\frac{d\delta_{x}{\cal S}_{2n}}{d\delta_{x}{\cal Q}_{V}}(y)\leq e^{-\nu_{V}(V)}~e^{W^{V}(x)+W_{U}(y)}$
@@
$e^{\nu_{V}(V)}~e^{-({\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}\leq\frac{d\delta_{y}{\cal S}_{2n+1}}{d\delta_{y}{\cal R}_{U}}(x)\leq e^{-\nu_{V}(V)}~e^{W^{V}(x)+W_{U}(y)}.$
@@
Using the Sinkhorn transition operators as expressed in (20) together with (131) and (132), we find that, for any $n\geq 1$,
${\cal S}_{2n}(x,dy)$ $=$ $\frac{{\cal Q}_{V}(x,dy)e^{(V-V_{2n})(y)}}{{\cal Q}(e^{-V_{2n}})(x)}\geq e^{\nu_{V}(V)}e^{-({\cal W}_{V,U}(x)+{\cal W}^{V,U}(y))}{\cal Q}_{V}(x,dy)$~
@@
${\cal S}_{2n+1}(y,dx)$ $=$ $\frac{{\cal R}_{U}(y,dx)~e^{(U-U_{2n})(x)}}{{\cal R}(e^{-U_{2n}})(y)}\geq e^{\nu_{V}(V)}~e^{-({\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}~{\cal R}_{U}(y,dx).$
Similarly, one can resort to (131) and (132) again to obtain the lower bounds
${\cal S}_{2n}(x,dy)$ $=$ $\frac{{\cal Q}_{V}(x,dy)e^{V-V_{2n}(y)}}{{\cal Q}(e^{-V_{2n}})(x)}\leq e^{-\nu_{V}(V)}~e^{W^{V}(x)}~{\cal Q}_{V}(x,dy)~e^{W_{U}(y)}$
@@
${\cal S}_{2n+1}(y,dx)$ $=$ $\frac{{\cal R}_{U}(y,dx)~e^{(U-U_{2n})(x)}}{{\cal R}(e^{-U_{2n}})(y)}\leq e^{-\nu_{V}(V)}~e^{W_{U}(y)}~{\cal R}_{U}(y,dx)~e^{W^{V}(x)}.$
@@
Assume that Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. Then, the functions $(\varphi,\psi)=(e^{\delta U},e^{\delta V})$ satisfy the Lyapunov inequalities
${\cal S}_{2n}\left(\psi\right)/\varphi\leq c_{\delta}~e^{-{\cal U}_{\delta}}~\in{\cal B}_{0}(\mathbb{X})\quad\mbox{and}\quad{\cal S}_{2n-1}\left(\varphi\right)/\psi\leq c_{\delta}~e^{-{\cal V}_{\delta}}~\in{\cal B}_{0}(\mathbb{Y})$ (137)
@@
${\cal S}_{2n}(x,dy)~\psi(y)/\varphi(x)$ $\leq$ $e^{-\nu_{V}(V)}~e^{-{\cal U}_{\delta}(x)}~{\cal Q}_{(1-\delta)V}(x,dy)~e^{W_{U}(y)},\quad\text{and}$
${\cal S}_{2n+1}(y,dx)~\varphi(x)/\psi(y)$ $\leq$ $e^{-\nu_{V}(V)}~e^{-{\cal V}_{\delta}(y)}~{\cal R}_{(1-\delta)U}(y,dx)~e^{W^{V}(x)},$
which (via (129)) imply the inequalities
${\cal S}_{2n}\left(\psi\right)/\varphi$ $\leq$ $e^{-\nu_{V}(V)}~e^{-{\cal U}_{\delta}}~e^{{\cal W}_{(1-\delta)V,U}},\quad\text{and}$
${\cal S}_{2n+1}\left(\varphi\right)/\psi$ $\leq$ $e^{-\nu_{V}(V)}~e^{-{\cal V}_{\delta}}~e^{{\cal W}^{(1-\delta)U,V}}.$
@@
$\|e^{{\cal W}_{(1-\delta)V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}=\|{\cal Q}_{(1-\delta)V}(e^{W_{U}})\|_{\mathbb{L}_{\infty}(\lambda)}\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\nu_{(1-\delta)V}\left(e^{W_{U}}\right)$
@@
$\|e^{{\cal W}^{(1-\delta)U,V}}\|_{\mathbb{L}_{\infty}(\nu)}=\|{\cal R}_{(1-\delta)U}(e^{W^{V}})\|_{\mathbb{L}_{\infty}(\nu)}\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\lambda_{(1-\delta)U}\left(e^{W^{V}}\right),$
where $\nu_{(1-\delta)V}\left(e^{W_{U}}\right)\vee\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)<\infty$ when Condition ${\cal H}_{\delta}(U,V)$ holds. This yields (137) with the constant
@@
By (130) we have
${\cal S}_{1}(y,dx)$ $=$ $\frac{{\cal R}(y,dx)~e^{-U(x)}}{{\cal R}(e^{-U})(y)}\leq e^{W_{U}(y)}~{\cal R}(y,dx)~e^{-U(x)}$
@@
$e^{-\delta V(y)}{\cal S}_{1}\left(e^{\delta U}\right)(y)\leq e^{-(\delta V(y)-W_{U}(y))}~\|e^{-W}\|~\lambda_{(1-\delta)U}\left(1\right)$
@@
${\cal S}_{1}\left(\varphi\right)/\psi\leq c_{\delta}~e^{-{\cal V}_{\delta}}$~
@@
D.4 Proof of Theorem 4.13
The Lyapunov inequalities (137) in Proposition D.9 ensure that the drift conditions stated in (60) hold with $n_{0}=1$ and $(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})$. Recall from the proof of Proposition D.5 that $e^{-\delta U}$ and $e^{-\delta V}$ are uniformly bounded for any $\delta\geq 0$. Since $\mathbb{X}$ and $\mathbb{Y}$ can be exhausted respectively by the compact sub-levels sets $C_{U}(r)$ and $C_{V}(r)$ of $U$ and $V$, there exists some $r_{1}>1$ such that $\lambda(C_{U}(r))\wedge\nu(C_{V}(r))>0$ for $r=r_{1}$ and thus for any $r\geq r_{1}$. This ensure there exists some $r_{0}>0$ such that for any $r\geq r_{0}$ we have
@@
${\cal S}_{2n}(x,dy)$ $\geq$ $e^{\nu_{V}(V)}~e^{-(W(x,y)+{\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}~\nu_{V}(dy)~1_{C_{\psi}(r)}(y)$
$\geq$ $e^{-w(r)}~\nu_{V}\left(C_{\psi}(r)\right)~{\cal Y}_{r}(dy)\quad\mbox{\rm with}\quad{\cal Y}_{r}(dy):=\frac{\nu_{V}(dy)~1_{C_{\psi}(r)}(y)}{\nu_{V}\left(C_{\psi}(r)\right)}$
@@
${\cal S}_{2n}(x,dy)\geq\varepsilon_{0}(r)~{\cal Y}_{r}(dy)\quad\mbox{\rm with}\quad\varepsilon_{0}(r):=w_{0}(r)~e^{-w(r)}.$
@@
${\cal S}_{2n+1}(y,dx)$ $\geq$ $e^{\nu_{V}(V)}~e^{-(W(x,y)+{\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}~\lambda_{U}(dx)~1_{C_{\varphi}(r)}(x)$
@@
$\varepsilon_{1}(r):=w_{1}(r)~e^{-w(r)}\quad\mbox{\rm and}\quad{\cal X}_{r}(dx):=\frac{\lambda_{U}(dx)~1_{C_{\varphi}(r)}(x)}{\lambda_{U}\left(C_{\varphi}(r)\right)}.$
The above estimates ensure that the local minorization condition (61) is met with $n_{0}=1$ and the function
@@
Hence, we conclude that Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied with $n_{0}=1$.
Using (22) and (125), for any $n\geq p\geq 1$ we verify that
$\frac{d\pi_{2n}}{d\nu_{V}}(y)-1=\left[(\delta_{y}-\nu_{V})\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\right)\right]\left(\frac{d\pi_{2p}}{d\nu_{V}}\right),$
@@
On the other hand, combining Proposition D.7 with the norm equivalence (34) and Theorem 4.8, we see that there exists some $\kappa>0$ and some constants $c_{\kappa,i}>0$, $i=1,2$, such that
@@
$\left\|\frac{d\pi_{2n}}{d\nu_{V}}-1\right\|_{\psi}\leq c_{\kappa,3}~\rho^{2n}~\quad\text{and, similarly,}\quad\left\|\frac{d\nu_{V}}{d\pi_{2n}}-1\right\|_{\psi}\leq c_{\kappa,4}~\rho^{2n},$ (138)
for some constants $c_{\kappa,i}$, $i=3,4$. By the same argument, using (22) we have, for any $n\geq p\geq 1$,
$\frac{d\pi_{2n+1}}{d\lambda_{U}}(x)-1=\left[(\delta_{x}-\lambda_{U})({\cal S}^{\circ}_{2n+1}\ldots{\cal S}^{\circ}_{2p+1})\right]\left(\frac{d\pi_{2p-1}}{d\lambda_{U}}\right)$
@@
$\left\|\frac{d\pi_{2n+1}}{d\lambda_{U}}-1\right\|_{\varphi}\vee\left\|\frac{d\lambda_{U}}{d\pi_{2n+1}}-1\right\|_{\varphi}\leq c_{\kappa,5}~\rho^{2n}$~ (139)
@@
$\|\log{f}\|_{\varphi}=\|\log{1/f}\|_{\varphi}\leq\|f-1\|_{\varphi}+\|1/f-1\|_{\varphi}$
The inequality above combined with (138) and (139) yields the estimates
@@
for some finite constant $c_{\kappa,6}$. This completes the proof of (71).
@@ Appendix E Some technical proofs
Proof of (123) and (124)
We first note, using (16) and (17), that
$\frac{d{\cal P}_{2n}}{d{\cal P}_{2n+1}}(x,y)=\frac{d\pi_{2n}}{d\nu_{V}}(y)=\exp{\left(V_{2n+1}(y)-V_{2n}(y)\right)}$ (140)
@@
$\frac{d{\cal P}_{2n+1}}{d{\cal P}_{2(n+1)}}(x,y)=\frac{d\pi_{2n+1}}{d\lambda_{U}}(x)=\exp{\left(U_{2(n+1)}(x)-U_{2n+1}(x)\right)}.$ (141)
@@
${\cal P}_{2(n+1)}(d(x,y))$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n+1}}(x)~{\cal P}_{2n}(d(x,y))~\frac{d\nu_{V}}{d\pi_{2n}}(y)\quad\text{and}$
${\cal P}_{2n+1}(d(x,y))$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n-1}}(x)~{\cal P}_{2n-1}(d(x,y))~\frac{d\nu_{V}}{d\pi_{2n}}(y)$
then, using (15), we obtain the formulae
${\cal S}_{2(n+1)}(x,dy)$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n+1}}(x)~{\cal S}_{2n}(x,dy)~\frac{d\nu_{V}}{d\pi_{2n}}(y)\quad\text{and}$
${\cal S}_{2n+1}(y,dx)$ $=$ $~\frac{d\nu_{V}}{d\pi_{2n}}(y)~{\cal S}_{2n-1}(y,dx)~\frac{d\lambda_{U}}{d\pi_{2n-1}}(x),$
@@
Proof of (57)
Assume Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied. Combining (53) with (22) we arrive at
$\nu_{V}(\psi)=\nu_{V}({\cal S}_{2n}^{\circ}(\psi))\leq\epsilon~\nu_{V}(\psi)+c\quad\text{which implies}\quad\nu_{V}(\psi)\leq\frac{c}{1-\epsilon}.$
@@
Using the Markov evolution equations in (23) together with (53) we find that, for any $n\geq n_{0}$,
$\pi_{2n}(\psi)$ $=$ $\pi_{2(n-1)}({\cal S}_{2n}^{\circ}(\psi))$
@@
$\pi_{2n+1}(\varphi)$ $=$ $\pi_{2n-1}({\cal S}_{2n+1}^{\circ}(\varphi))$
@@
${\cal V}_{\delta}=\delta V-W_{U}\in{\cal B}_{\infty}(\mathbb{R}^{d})\quad\mbox{\rm and}\quad{\cal U}_{\delta}=\delta U-W^{V}\in{\cal B}_{\infty}(\mathbb{R}^{d})$
@@
$W_{U}=\sum_{i\in E_{0}}~\xi_{0}(i)~W_{U_{i}}\quad\mbox{\rm and}\quad W^{V}=\sum_{j\in E_{1}}~\xi_{1}(j)~W^{V_{j}}$ (142)
@@
and this ensures that ${\cal U}_{\delta}=\delta U-W^{V}$ have compact sub-level sets. In the same vein, one can also prove that ${\cal V}_{\delta}=\delta V-W_{U}$ have compact sub-level sets.
@@
$\lambda_{(1-\delta)U}(dx)$ $=$ $\left(\sum_{i\in E_{0}}\xi_{0}(i)~e^{-U_{i}(x)}\right)^{1-\delta}~\lambda(dx)$ (143)
@@
$\lambda_{(1-\delta)U_{i}}\left(\prod_{j\in E_{1}}~e^{\xi_{1}(j)~W^{V_{j}}}\right)\leq\prod_{j\in E_{1}}~\left(\lambda_{(1-\delta)U_{i}}\left(e^{W^{V_{j}}}\right)\right)^{\xi_{1}(j)}.$ (144)
Combining (142), (143) and (144) we obtain the estimates

### v3

@@ D.3. Lyapunov inequalities
$e^{\nu_{V}(V)}~e^{-({\mathcal{W}}_{V,U}(x)+{\mathcal{W}}^{U,V}(y))}\leq\frac{d\delta_{x}{\mathcal{S}}_{2n}}{d\delta_{x}{\mathcal{Q}}_{V}}(y)\leq e^{-\nu_{V}(V)}~e^{W^{V}(x)+W_{U}(y)}$
@@
$e^{\nu_{V}(V)}~e^{-({\mathcal{W}}_{V,U}(x)+{\mathcal{W}}^{U,V}(y))}\leq\frac{d\delta_{y}{\mathcal{S}}_{2n+1}}{d\delta_{y}{\mathcal{R}}_{U}}(x)\leq e^{-\nu_{V}(V)}~e^{W^{V}(x)+W_{U}(y)}.$
@@
Using the Sinkhorn transition operators as expressed in (2.6) together with (D.9) and (D.10), we find that, for any $n\geq 1$,
${\mathcal{S}}_{2n}(x,dy)$ $=$ $\frac{{\mathcal{Q}}_{V}(x,dy)e^{(V-V_{2n})(y)}}{{\mathcal{Q}}(e^{-V_{2n}})(x)}\geq e^{\nu_{V}(V)}e^{-({\mathcal{W}}_{V,U}(x)+{\mathcal{W}}^{V,U}(y))}{\mathcal{Q}}_{V}(x,dy)$~
@@
${\mathcal{S}}_{2n+1}(y,dx)$ $=$ $\frac{{\mathcal{R}}_{U}(y,dx)~e^{(U-U_{2n})(x)}}{{\mathcal{R}}(e^{-U_{2n}})(y)}\geq e^{\nu_{V}(V)}~e^{-({\mathcal{W}}_{V,U}(x)+{\mathcal{W}}^{U,V}(y))}~{\mathcal{R}}_{U}(y,dx).$
Similarly, one can resort to (D.9) and (D.10) again to obtain the lower bounds
${\mathcal{S}}_{2n}(x,dy)$ $=$ $\frac{{\mathcal{Q}}_{V}(x,dy)e^{V-V_{2n}(y)}}{{\mathcal{Q}}(e^{-V_{2n}})(x)}\leq e^{-\nu_{V}(V)}~e^{W^{V}(x)}~{\mathcal{Q}}_{V}(x,dy)~e^{W_{U}(y)}$
@@
${\mathcal{S}}_{2n+1}(y,dx)$ $=$ $\frac{{\mathcal{R}}_{U}(y,dx)~e^{(U-U_{2n})(x)}}{{\mathcal{R}}(e^{-U_{2n}})(y)}\leq e^{-\nu_{V}(V)}~e^{W_{U}(y)}~{\mathcal{R}}_{U}(y,dx)~e^{W^{V}(x)}.$
@@
Assume that Condition ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. Then, the functions $(\varphi,\psi)=(e^{\delta U},e^{\delta V})$ satisfy the Lyapunov inequalities
(D.15) ${\mathcal{S}}_{2n}\left(\psi\right)/\varphi\leq c_{\delta}~e^{-{\mathcal{U}}_{\delta}}~\in{\mathcal{B}}_{0}(\mathbb{X})\quad\mbox{and}\quad{\mathcal{S}}_{2n-1}\left(\varphi\right)/\psi\leq c_{\delta}~e^{-{\mathcal{V}}_{\delta}}~\in{\mathcal{B}}_{0}(\mathbb{Y})$
@@
${\mathcal{S}}_{2n}(x,dy)~\psi(y)/\varphi(x)$ $\leq$ $e^{-\nu_{V}(V)}~e^{-{\mathcal{U}}_{\delta}(x)}~{\mathcal{Q}}_{(1-\delta)V}(x,dy)~e^{W_{U}(y)},\quad\text{and}$
${\mathcal{S}}_{2n+1}(y,dx)~\varphi(x)/\psi(y)$ $\leq$ $e^{-\nu_{V}(V)}~e^{-{\mathcal{V}}_{\delta}(y)}~{\mathcal{R}}_{(1-\delta)U}(y,dx)~e^{W^{V}(x)},$
which (via (D.7)) imply the inequalities
${\mathcal{S}}_{2n}\left(\psi\right)/\varphi$ $\leq$ $e^{-\nu_{V}(V)}~e^{-{\mathcal{U}}_{\delta}}~e^{{\mathcal{W}}_{(1-\delta)V,U}},\quad\text{and}$
${\mathcal{S}}_{2n+1}\left(\varphi\right)/\psi$ $\leq$ $e^{-\nu_{V}(V)}~e^{-{\mathcal{V}}_{\delta}}~e^{{\mathcal{W}}^{(1-\delta)U,V}}.$
@@
$\|e^{{\mathcal{W}}_{(1-\delta)V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}=\|{\mathcal{Q}}_{(1-\delta)V}(e^{W_{U}})\|_{\mathbb{L}_{\infty}(\lambda)}\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\nu_{(1-\delta)V}\left(e^{W_{U}}\right)$
@@
$\|e^{{\mathcal{W}}^{(1-\delta)U,V}}\|_{\mathbb{L}_{\infty}(\nu)}=\|{\mathcal{R}}_{(1-\delta)U}(e^{W^{V}})\|_{\mathbb{L}_{\infty}(\nu)}\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\lambda_{(1-\delta)U}\left(e^{W^{V}}\right),$
where $\nu_{(1-\delta)V}\left(e^{W_{U}}\right)\vee\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)<\infty$ when Condition ${\mathcal{H}}_{\delta}(U,V)$ holds. This yields (D.15) with the constant
@@
By (D.8) we have
${\mathcal{S}}_{1}(y,dx)$ $=$ $\frac{{\mathcal{R}}(y,dx)~e^{-U(x)}}{{\mathcal{R}}(e^{-U})(y)}\leq e^{W_{U}(y)}~{\mathcal{R}}(y,dx)~e^{-U(x)}$
@@
$e^{-\delta V(y)}{\mathcal{S}}_{1}\left(e^{\delta U}\right)(y)\leq e^{-(\delta V(y)-W_{U}(y))}~\|e^{-W}\|~\lambda_{(1-\delta)U}\left(1\right)$
@@
${\mathcal{S}}_{1}\left(\varphi\right)/\psi\leq c_{\delta}~e^{-{\mathcal{V}}_{\delta}}$~
@@
D.4. Proof of Theorem 4.8
By (C.7), the Lyapunov inequalities (D.15) in Proposition D.9 ensure that the drift conditions stated in (4.13) hold with $n_{0}=1$ and $(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})$. Recall from the proof of Proposition D.5 that $e^{-\delta U}$ and $e^{-\delta V}$ are uniformly bounded for any $\delta\geq 0$. Since $\mathbb{X}$ and $\mathbb{Y}$ can be exhausted respectively by the compact sub-levels sets $C_{U}(r)$ and $C_{V}(r)$ of $U$ and $V$, there exists some $r_{1}>1$ such that $\lambda(C_{U}(r))\wedge\nu(C_{V}(r))>0$ for $r=r_{1}$ and thus for any $r\geq r_{1}$. This ensures there exists some $r_{0}>0$ such that for any $r\geq r_{0}$ we have
@@
${\mathcal{S}}_{2n}(x,dy)$ $\geq$ $e^{\nu_{V}(V)}~e^{-(W(x,y)+{\mathcal{W}}_{V,U}(x)+{\mathcal{W}}^{U,V}(y))}~\nu_{V}(dy)~1_{C_{\psi}(r)}(y)$
$\geq$ $e^{-w(r)}~\nu_{V}\left(C_{\psi}(r)\right)~{\mathcal{Y}}_{r}(dy)\quad\mbox{\rm with}\quad{\mathcal{Y}}_{r}(dy):=\frac{\nu_{V}(dy)~1_{C_{\psi}(r)}(y)}{\nu_{V}\left(C_{\psi}(r)\right)}$
@@
${\mathcal{S}}_{2n}(x,dy)\geq\varepsilon_{0}(r)~{\mathcal{Y}}_{r}(dy)\quad\mbox{\rm with}\quad\varepsilon_{0}(r):=w_{0}(r)~e^{-w(r)}.$
@@
${\mathcal{S}}_{2n+1}(y,dx)$ $\geq$ $e^{\nu_{V}(V)}~e^{-(W(x,y)+{\mathcal{W}}_{V,U}(x)+{\mathcal{W}}^{U,V}(y))}~\lambda_{U}(dx)~1_{C_{\varphi}(r)}(x)$
@@
$\varepsilon_{1}(r):=w_{1}(r)~e^{-w(r)}\quad\mbox{\rm and}\quad{\mathcal{X}}_{r}(dx):=\frac{\lambda_{U}(dx)~1_{C_{\varphi}(r)}(x)}{\lambda_{U}\left(C_{\varphi}(r)\right)}.$
The above estimates ensure that the local minorization condition (4.14) is met with $n_{0}=1$ and the function
@@
Hence, we conclude that Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ is satisfied with $n_{0}=1$.
Using (2.8) and (D.3), for any $n\geq p\geq 1$ we verify that
$\frac{d\pi_{2n}}{d\nu_{V}}(y)-1=\left[(\delta_{y}-\nu_{V})\left({\mathcal{S}}^{\circ}_{2n}\ldots{\mathcal{S}}^{\circ}_{2(p+1)}\right)\right]\left(\frac{d\pi_{2p}}{d\nu_{V}}\right),$
@@
On the other hand, combining Proposition D.7 with the norm equivalence (3.4) and Theorem 4.6, we see that there exists some $\kappa>0$ and some constants $c_{\kappa,i}>0$, $i=1,2$, such that
@@
(D.16) $\left\|\frac{d\pi_{2n}}{d\nu_{V}}-1\right\|_{\psi}\leq c_{\kappa,3}~\rho^{2n}~\quad\text{and, similarly,}\quad\left\|\frac{d\nu_{V}}{d\pi_{2n}}-1\right\|_{\psi}\leq c_{\kappa,4}~\rho^{2n},$
for some constants $c_{\kappa,i}$, $i=3,4$. By the same argument, using (2.8) we have, for any $n\geq p\geq 1$,
$\frac{d\pi_{2n+1}}{d\lambda_{U}}(x)-1=\left[(\delta_{x}-\lambda_{U})({\mathcal{S}}^{\circ}_{2n+1}\ldots{\mathcal{S}}^{\circ}_{2p+1})\right]\left(\frac{d\pi_{2p-1}}{d\lambda_{U}}\right)$
@@
(D.17) $\left\|\frac{d\pi_{2n+1}}{d\lambda_{U}}-1\right\|_{\varphi}\vee\left\|\frac{d\lambda_{U}}{d\pi_{2n+1}}-1\right\|_{\varphi}\leq c_{\kappa,5}~\rho^{2n}$~
@@
$\|\log{f}\|_{\varphi}=\|\log{1/f}\|_{\varphi}\leq\|f-1\|_{\varphi}+\|1/f-1\|_{\varphi}.$
The inequality above combined with (D.16) and (D.17) yields the estimates
@@
for some finite constant $c_{\kappa,6}$. This completes the proof of (4.24).
@@ Appendix E Some technical proofs
Proof of (D.1) and (D.2)
We first note, using (2.2) and (2.3), that
(E.1) $\frac{d{\mathcal{P}}_{2n}}{d{\mathcal{P}}_{2n+1}}(x,y)=\frac{d\pi_{2n}}{d\nu_{V}}(y)=\exp{\left(V_{2n+1}(y)-V_{2n}(y)\right)}$
@@
(E.2) $\frac{d{\mathcal{P}}_{2n+1}}{d{\mathcal{P}}_{2(n+1)}}(x,y)=\frac{d\pi_{2n+1}}{d\lambda_{U}}(x)=\exp{\left(U_{2(n+1)}(x)-U_{2n+1}(x)\right)}.$
@@
${\mathcal{P}}_{2(n+1)}(d(x,y))$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n+1}}(x)~{\mathcal{P}}_{2n}(d(x,y))~\frac{d\nu_{V}}{d\pi_{2n}}(y)\quad\text{and}$
${\mathcal{P}}_{2n+1}(d(x,y))$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n-1}}(x)~{\mathcal{P}}_{2n-1}(d(x,y))~\frac{d\nu_{V}}{d\pi_{2n}}(y)$
then, using (2.1), we obtain the formulae
${\mathcal{S}}_{2(n+1)}(x,dy)$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n+1}}(x)~{\mathcal{S}}_{2n}(x,dy)~\frac{d\nu_{V}}{d\pi_{2n}}(y)\quad\text{and}$
${\mathcal{S}}_{2n+1}(y,dx)$ $=$ $~\frac{d\nu_{V}}{d\pi_{2n}}(y)~{\mathcal{S}}_{2n-1}(y,dx)~\frac{d\lambda_{U}}{d\pi_{2n-1}}(x),$
@@
Proof of (4.10)
Assume Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ is satisfied. Combining (4.6) with (2.8) we arrive at
$\nu_{V}(\psi)=\nu_{V}({\mathcal{S}}_{2n}^{\circ}(\psi))\leq\epsilon~\nu_{V}(\psi)+c\quad\text{which implies}\quad\nu_{V}(\psi)\leq\frac{c}{1-\epsilon}.$
@@
Using the Markov evolution equations in (2.9) together with (4.6) we find that, for any $n\geq n_{0}$,
$\pi_{2n}(\psi)$ $=$ $\pi_{2(n-1)}({\mathcal{S}}_{2n}^{\circ}(\psi))$
@@
$\pi_{2n+1}(\varphi)$ $=$ $\pi_{2n-1}({\mathcal{S}}_{2n+1}^{\circ}(\varphi))$
@@
${\mathcal{V}}_{\delta}=\delta V-W_{U}\in{\mathcal{B}}_{\infty}(\mathbb{R}^{d})\quad\mbox{\rm and}\quad{\mathcal{U}}_{\delta}=\delta U-W^{V}\in{\mathcal{B}}_{\infty}(\mathbb{R}^{d})$
@@
(E.3) $W_{U}=\sum_{i\in E_{0}}~\xi_{0}(i)~W_{U_{i}}\quad\mbox{\rm and}\quad W^{V}=\sum_{j\in E_{1}}~\xi_{1}(j)~W^{V_{j}}$
@@
and this ensures that ${\mathcal{U}}_{\delta}=\delta U-W^{V}$ have compact sub-level sets. In the same vein, one can also prove that ${\mathcal{V}}_{\delta}=\delta V-W_{U}$ have compact sub-level sets.
@@
(E.4) $\lambda_{(1-\delta)U}(dx)$ $=$ $\left(\sum_{i\in E_{0}}\xi_{0}(i)~e^{-U_{i}(x)}\right)^{1-\delta}~\lambda(dx)$
@@
(E.5) $\lambda_{(1-\delta)U_{i}}\left(\prod_{j\in E_{1}}~e^{\xi_{1}(j)~W^{V_{j}}}\right)\leq\prod_{j\in E_{1}}~\left(\lambda_{(1-\delta)U_{i}}\left(e^{W^{V_{j}}}\right)\right)^{\xi_{1}(j)}.$
Combining (E.3), (E.4) and (E.5) we obtain the estimates